# 1NLPR, MAIS, CASIA 2UCAS 3PKU 4WHU 5ByteDance

### v2

@@ 4 Experiments
| Osprey-7B [60] | 39.9 | 45.9 | 47.1 | 30.0 | 48.6 | 23.7 | |||||||||||||||||
| Ferret-7B [58] | 48.8 | 35.2 | 44.7 | 41.9 | 70.4 | 74.6 | |||||||||||||||||
| VideoRefer-7B† [61] | 71.9 | 75.4 | 68.6 | 59.3 | 89.4 | 78.1 | |||||||||||||||||
@@
Qualitative Results. We provide qualitative comparisons between our GAR-8B with DAM-3B [22] on detailed localized captioning on DLC-Bench [22] in Figure 5. As demonstrated in the figure, our GAR-8B is more capable of generating precise descriptions, especially when the category of the given prompt can be determined only when understanding sufficient global contexts. More comparisons can be found in § D.
@@ 5 Conclusion
This paper introduces Grasp Any Region (GAR), a family of MLLMs for region understanding, and GAR-Bench, a systematic evaluation framework that not only provides a more accurate evaluation of single-region comprehension, but also for multi-prompt interaction and advanced compositional reasoning. On detailed captioning benchmarks [22, 58, 24], GAR demonstrates superior performance over DAM [22]. More importantly, our GAR achieves advanced comprehension capability in modeling interactions between multiple prompts. Specifically, on GAR-Bench-VQA, GAR-1B even surpasses InternVL3-78B [66]. On VideoRefer-Bench [61], our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B [61]. We hope our work inspires the community to develop MLLMs that can perceive, interrogate, and understand the dense visual world more effectively.
@@
Our research is grounded in ethical practices, with particular attention paid to the responsible use of data. All datasets employed in this study are publicly available and well-established within the computer vision community. Specifically, our training data includes ImageNet-21K [11] and the PSG [55] dataset (with image sources from COCO [23], while our benchmarking was conducted on FSC-147 [37], RGBD-Mirror [30], and SA-1B [17]. Our use of this data is in accordance with their provided licenses and intended academic purpose.
@@
We are committed to ensuring the reproducibility of the research presented in this paper. To this end, comprehensive implementation details for our models and experiments are provided in § C, including the training procedures and all hyperparameters used. Furthermore, upon acceptance of this paper, all source code, datasets, and trained model checkpoints will be made publicly available.
@@ Appendix A Overview
In § B, we introduce details of our GAR-Bench, including the annotation pipeline and statistics.
In § C, we provide more implementation details as well as experimental results. Detailed ablations of each component can be found in this section.
In § D, we provide qualitative results on both detailed image captioning and understanding, and localized video captioning and understanding.
In § E, we discuss potential limitations and analyze failure cases.
In § G, we provide all prompts we utilized to construct our dataset.
Finally in § H, we discuss the use of LLMs in preparing this paper.
@@ B.1 Annotation Pipeline
Image Selection. To ensure the relevance and challenge of our sub-tasks, we begin by carefully curating source images from existing datasets known to contain specific visual patterns. For the “relation” tasks, we source images from the Panoptic Scene Graph (PSG) dataset [55], which is rich in complex scene graphs and explicit object relationships, providing a natural foundation for multi-prompt interaction queries. For the “non-entity recognition” task, we utilize the RGBD-Mirror dataset [30], as it specifically contains scenes with mirrors and reflections, allowing us to create unambiguous test cases for distinguishing real objects from illusory ones. For the “position” task, we select images from the FSC-147 dataset [37], which features images with numerous countable objects often arranged in grid-like patterns, making it ideal for evaluating spatial and ordinal reasoning. Other images are from SA-1B [17].
Mask Labeling. Following image selection, we generate high-quality segmentation masks for all potential objects of interest. This stage is similar to [21], which decomposes complex scenes into different objects, while not containing numerous meaningless, trivial objects like those in the SA-1B [17] dataset.
@@
Automated Attribute Generation. For the “perception” protocol of GAR-Bench-VQA, we leverage the advanced capabilities of Gemini-2.5-Pro [10]. For each selected and verified object mask, we prompt the model to generate a list of its basic perceptual attributes, including its primary color, shape, material, and any discernible texture or pattern.
Quality Control and Formatting. The raw, annotated data then underwent a meticulous, multi-stage quality control process. First, human experts review all machine-generated attributes from the previous step to verify their factual correctness and filter out any ambiguous or inaccurate labels. Following this verification, the experts transform the raw annotations into the final benchmark formats. For all VQA tasks, they rewrite the question-answer pairs into a standardized multiple-choice format, ensuring consistent and objective evaluation. For the captioning task, the ground-truth data was structured for compatibility with LLM-as-a-Judge evaluation protocols similar to [22].
Difficulty Filtering. As a final quality assurance measure, we implement a difficulty filtering process to ensure the benchmark remains challenging for even the most advanced models. Specifically, any question answered correctly by all four state-of-the-art non-thinking MLLMs, i.e., Qwen2.5-VL-72B [1], InternVL3-78B [66], GPT-4o [31], and Gemini-2.5-Flash [9], was excluded from the final benchmark.
@@
| Global | Local | GAR-Bench | DLC-Bench† | ||||
| Caption | VQA | Avg. | Pos. | Neg. | |||
| Base Model: Qwen2.5-VL-3B [1] | |||||||
@@
| Base Model: InternVL3-2B [66] | |||||||
@@ Appendix C More Experiments
Implementation Details. We adopt PerceptionLM series [7] as our base model, as it demonstrates strong perception capabilities among several open-source MLLMs. We perform supervised fine-tuning of the model on our GAR-2.5M using Xtuner [8] with the AdamW optimizer [29] with a global batch size of 64 and a learning rate of 1e-5 with a cosine decay [28].
Comparison Baselines. We mainly compare our GAR with both general MLLMs, including state-of-the-art private models [31, 33, 10], and representative public models [1, 66, 26], and region-level MLLMs, including GLaMM [38], GPT4RoI [64], Osprey [60], Shikra [3], Ferret [58], RegionGPT [14], OMG-LLaVA [65], VP-SPHINX [24], Sa2VA [59], DAM [22], and PAM [25]. We transform masks to boxes for box-level MLLMs, e.g., [64, 3, 58, 25], as our GAR-Bench provides segmentation masks by default. On video benchmarks, we further compare with LLaVA-OneVision [20], Qwen2-VL [49], InternVL2 [6], Elysium [44], Artemis [34], and VideoRefer [61].
@@
In Table 9, we further extend our ablations on model architectures to more base models, including Qwen2.5-VL-3B [1] and InternVL3-2B [66]. As demonstrated in the table, our proposed RoI-aligned feature replay consistently brings significant improvements over different base models.
@@
Ablations on Data Pipeline. We study the effectiveness of our data in Table 11. Starting from the seed dataset, i.e., Describe-Anything-1.5M [22], we first add our Fine-Grained Dataset-456K, and then add our Relation Dataset-414K. By introducing our Fine-Grained Dataset-456K, our model is able to produce more accurate recognition, leading to an improvement of +3.1 on DLC-Bench [22]. By further combining our proposed Relation Dataset-414K, the model is finally equipped with compositional reasoning capabilities with multiple prompts at this time, resulting in significant improvements on our GAR-Bench.
@@
Performances on General Multimodal Benchmarks. We compare our GAR-8B with other region-level models, i.e., DAM-3B [22] and PAM-3B [25], on general vision-centric multimodal benchmarks, including V* [51], MMVP [43], RealWorldQA [53], and MMStar [4]. As illustrated in Table 11, our GAR-8B outperforms them by a large margin.
@@ D.1 Qualitative Results on GAR-Bench
“Relation” of GAR-Bench-VQA. In Figure 7, we provide qualitative comparisons on the “relation” protocol of our GAR-Bench-VQA, including two failure cases (the last row). As demonstrated in the figure, GAR-8B manages to not only effectively model relationships but also leverage crucial local details for choosing the best answer. For instance, in the right example of the middle row, the person (<Prompt0>) is actually not reading the book (<Prompt1>), since she is looking at the camera. Our GAR-8B manages to recognize such details and thus select “<Prompt0> is holding <Prompt1>” instead of “reading”, while both Gemini-2.5-Pro [10] and o3 [33] fail. However, as illustrated in the last two examples in Figure 7, current models still sometimes struggle to understand complex relationships with more than two objects. Constructing such complicated training data and keeping the correctness of relation annotations could be a potential solution.
“Non-Entity Recognition” of GAR-Bench-VQA. In Figure 8, we provide qualitative comparisons on the “non-entity recognition” protocol of our GAR-Bench-VQA, including two failure cases (the last row). As demonstrated in the figure, GAR-8B is able to correctly recognize objects in the mirror without any depth prior, thanks to its encoded global contexts. However, as demonstrated in the right case in the last row, current models still struggle to distinguish whether the reflection actually comes from the mirror (<Prompt2>) or other reflective surfaces (<Prompt0> and <Prompt1>).

### v3

@@ 4 Experiments
| Osprey-7B | 39.9 | 45.9 | 47.1 | 30.0 | 48.6 | 23.7 | |||||||||||||||||
| Ferret-7B | 48.8 | 35.2 | 44.7 | 41.9 | 70.4 | 74.6 | |||||||||||||||||
| VideoRefer-7B† | 71.9 | 75.4 | 68.6 | 59.3 | 89.4 | 78.1 | |||||||||||||||||
@@
Qualitative Results. We provide qualitative comparisons between our GAR-8B with DAM-3B (Lian et al., 2025) on detailed localized captioning on DLC-Bench (Lian et al., 2025) in Figure 5. As demonstrated in the figure, our GAR-8B is more capable of generating precise descriptions, especially when the category of the given prompt can be determined only when understanding sufficient global contexts. More comparisons can be found in Appendix D.
@@ 5 Conclusion
This paper introduces Grasp Any Region (GAR), a family of MLLMs for region understanding, and GAR-Bench, a systematic evaluation framework that not only provides a more accurate evaluation of single-region comprehension, but also for multi-prompt interaction and advanced compositional reasoning. On detailed captioning benchmarks (Lian et al., 2025; You et al., 2023; Lin et al., 2025a), GAR demonstrates superior performance over DAM (Lian et al., 2025). More importantly, our GAR achieves advanced comprehension capability in modeling interactions between multiple prompts. Specifically, on GAR-Bench-VQA, GAR-1B even surpasses InternVL3-78B (Zhu et al., 2025). On VideoRefer-Bench (Yuan et al., 2025b), our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B (Yuan et al., 2025b). We hope our work inspires the community to develop MLLMs that can perceive, interrogate, and understand the dense visual world more effectively.
Acknowledgements
This work was supported by the Beijing Natural Science Foundation (No. L257015) and the National Natural Science Foundation of China (No. 62320106010).
@@ Acknowledgements
Our research is grounded in ethical practices, with particular attention paid to the responsible use of data. All datasets employed in this study are publicly available and well-established within the computer vision community. Specifically, our training data includes ImageNet-21K (Deng et al., 2009) and the PSG (Yang et al., 2022) dataset (with image sources from COCO (Lin et al., 2014), while our benchmarking was conducted on FSC-147 (Ranjan et al., 2021), RGBD-Mirror (Mei et al., 2021), and SA-1B (Kirillov et al., 2023). Our use of this data is in accordance with their provided licenses and intended academic purpose.
@@
We are committed to ensuring the reproducibility of the research presented in this paper. To this end, comprehensive implementation details for our models and experiments are provided in Appendix C, including the training procedures and all hyperparameters used. Furthermore, upon acceptance of this paper, all source code, datasets, and trained model checkpoints will be made publicly available.
@@ Appendix A Overview
In Appendix B, we introduce details of our GAR-Bench, including the annotation pipeline and statistics.
In Appendix C, we provide more implementation details as well as experimental results. Detailed ablations of each component can be found in this section.
In Appendix D, we provide qualitative results on both detailed image captioning and understanding, and localized video captioning and understanding.
In Appendix E, we discuss potential limitations and analyze failure cases.
In Appendix G, we provide all prompts we utilized to construct our dataset.
Finally in Appendix H, we discuss the use of LLMs in preparing this paper.
@@ B.1 Annotation Pipeline
Image Selection. To ensure the relevance and challenge of our sub-tasks, we begin by carefully curating source images from existing datasets known to contain specific visual patterns. For the “relation” tasks, we source images from the Panoptic Scene Graph (PSG) dataset (Yang et al., 2022), which is rich in complex scene graphs and explicit object relationships, providing a natural foundation for multi-prompt interaction queries. For the “non-entity recognition” task, we utilize the RGBD-Mirror dataset (Mei et al., 2021), as it specifically contains scenes with mirrors and reflections, allowing us to create unambiguous test cases for distinguishing real objects from illusory ones. For the “position” task, we select images from the FSC-147 dataset (Ranjan et al., 2021), which features images with numerous countable objects often arranged in grid-like patterns, making it ideal for evaluating spatial and ordinal reasoning. Other images are from SA-1B (Kirillov et al., 2023).
Mask Labeling. Following image selection, we generate high-quality segmentation masks for all potential objects of interest. This stage is similar to Li et al. (2025), which decomposes complex scenes into different objects, while not containing numerous meaningless, trivial objects like those in the SA-1B (Kirillov et al., 2023) dataset.
@@
Automated Attribute Generation. For the “perception” protocol of GAR-Bench-VQA, we leverage the advanced capabilities of Gemini-2.5-Pro (DeepMind, 2025b). For each selected and verified object mask, we prompt the model to generate a list of its basic perceptual attributes, including its primary color, shape, material, and any discernible texture or pattern.
Quality Control and Formatting. The raw, annotated data then underwent a meticulous, multi-stage quality control process. First, human experts review all machine-generated attributes from the previous step to verify their factual correctness and filter out any ambiguous or inaccurate labels. Following this verification, the experts transform the raw annotations into the final benchmark formats. For all VQA tasks, they rewrite the question-answer pairs into a standardized multiple-choice format, ensuring consistent and objective evaluation. For the captioning task, the ground-truth data was structured for compatibility with LLM-as-a-Judge evaluation protocols similar to Lian et al. (2025).
Difficulty Filtering. As a final quality assurance measure, we implement a difficulty filtering process to ensure the benchmark remains challenging for even the most advanced models. Specifically, any question answered correctly by all four state-of-the-art non-thinking MLLMs, i.e., Qwen2.5-VL-72B (Bai et al., 2025), InternVL3-78B (Zhu et al., 2025), GPT-4o (OpenAI, 2024a), and Gemini-2.5-Flash (DeepMind, 2025a), was excluded from the final benchmark.
@@
| Global | Local | GAR-Bench | DLC-Bench† | ||||
| Caption | VQA | Avg. | Pos. | Neg. | |||
| Base Model: Qwen2.5-VL-3B | |||||||
[…]
| Base Model: InternVL3-2B | |||||||
@@
| Method | V* | MMVP | RealWorldQA | MMStar |
| DAM-3B (Lian et al., 2025) | 45.0 | 60.7 | 54.3 | 39.7 |
| PAM-3B (Lin et al., 2025b) | 1.4 | 4.3 | 1.7 | 2.7 |
| Base Model: PerceptionLM-8B | ||||
| PercetionLM-8B (Cho et al., 2025) | 69.1 | 76.0 | 75.0 | 57.1 |
| GAR-8B | 59.2 | 78.0 | 58.7 | 43.9 |
| GAR-8B (w/ 600K General Data (Li et al., 2024)) | 62.3 | 79.7 | 61.8 | 51.6 |
@@ Appendix C More Experiments
Implementation Details. We adopt PerceptionLM series (Cho et al., 2025) as our base model, as it demonstrates strong perception capabilities among several open-source MLLMs. We perform supervised fine-tuning of the model on our GAR-2.5M using Xtuner (Contributors, 2023) with the AdamW optimizer (Loshchilov & Hutter, 2017) with a global batch size of 64 and a learning rate of 1e-5 with a cosine decay (Loshchilov & Hutter, 2016).
Comparison Baselines. We mainly compare our GAR with both general MLLMs, including state-of-the-art private models (OpenAI, 2024a; 2025; DeepMind, 2025b), and representative public models (Bai et al., 2025; Zhu et al., 2025; Liu et al., 2023), and region-level MLLMs, including GLaMM (Rasheed et al., 2024), GPT4RoI (Zhang et al., 2024a), Osprey (Yuan et al., 2024), Shikra (Chen et al., 2023), Ferret (You et al., 2023), RegionGPT (Guo et al., 2024), OMG-LLaVA (Zhang et al., 2024b), VP-SPHINX (Lin et al., 2025a), Sa2VA (Yuan et al., 2025a), DAM (Lian et al., 2025), and PAM (Lin et al., 2025b). We transform masks to boxes for box-level MLLMs, e.g., (Zhang et al., 2024a; Chen et al., 2023; You et al., 2023; Lin et al., 2025b), as our GAR-Bench provides segmentation masks by default. On video benchmarks, we further compare with LLaVA-OneVision (Li et al., 2024), Qwen2-VL (Wang et al., 2024b), InternVL2 (Chen et al., 2024c), Elysium (Wang et al., 2024a), Artemis (Qiu et al., 2024), and VideoRefer (Yuan et al., 2025b).
@@
In Table 9, we further extend our ablations on model architectures to more base models, including Qwen2.5-VL-3B (Bai et al., 2025) and InternVL3-2B (Zhu et al., 2025). As demonstrated in the table, our proposed RoI-aligned feature replay consistently brings significant improvements over different base models.
@@
Ablations on Data Pipeline. We study the effectiveness of our data in Table 11. Starting from the seed dataset, i.e., Describe-Anything-1.5M (Lian et al., 2025), we first add our Fine-Grained Dataset-456K, and then add our Relation Dataset-414K. By introducing our Fine-Grained Dataset-456K, our model is able to produce more accurate recognition, leading to an improvement of +3.1 on DLC-Bench (Lian et al., 2025). By further combining our proposed Relation Dataset-414K, the model is finally equipped with compositional reasoning capabilities with multiple prompts at this time, resulting in significant improvements on our GAR-Bench.
@@
Performances on General Multimodal Benchmarks. We compare our GAR-8B with other region-level models, i.e., DAM-3B (Lian et al., 2025) and PAM-3B (Lin et al., 2025b), on general vision-centric multimodal benchmarks, including V* (Wu & Xie, 2024), MMVP (Tong et al., 2024b), RealWorldQA (xAI, 2024), and MMStar (Chen et al., 2024a). As illustrated in Table 11, our GAR-8B outperforms them by a large margin.
| Model | Full (424) | 1/2 (212) | 1/4 (106) | |||
| Overall | Rank | Overall | Rank | Overall | Rank | |
| PAM-3B | 2.4 | 9 | 4.3 | 9 | 0.9 | 9 |
| VP-SPHINX-13B | 37.5 | 8 | 40.0 | 8 | 33.3 | 8 |
| DAM-3B | 38.2 | 7 | 48.6 | 7 | 41.9 | 7 |
| GAR-1B | 50.6 | 6 | 51.4 | 6 | 49.5 | 5 |
| Qwen2.5-VL-32B | 50.9 | 5 | 52.4 | 5 | 48.6 | 6 |
| GPT-4o | 53.5 | 4 | 56.7 | 4 | 57.1 | 4 |
| GAR-8B | 59.9 | 3 | 60.0 | 3 | 63.8 | 1 |
| o3 | 61.3 | 2 | 63.3 | 1 | 58.1 | 3 |
| Gemini-2.5-Pro | 64.2 | 1 | 61.0 | 2 | 60.0 | 2 |
| Model | Full (214) | 1/2 (107) | 1/4 (53) | |||
| Overall | Rank | Overall | Rank | Overall | Rank | |
| DAM-3B | 13.1 | 9 | 13.8 | 9 | 14.0 | 9 |
| PAM-3B | 21.1 | 8 | 18.8 | 8 | 20.1 | 8 |
| VP-SPHINX-13B | 32.3 | 7 | 29.7 | 7 | 20.0 | 7 |
| Qwen2.5-VL-32B | 36.8 | 6 | 32.7 | 6 | 26.1 | 6 |
| GPT-4o | 51.5 | 5 | 45.5 | 5 | 52.1 | 4 |
| o3 | 56.9 | 4 | 50.6 | 4 | 50.3 | 5 |
| GAR-1B | 57.5 | 3 | 51.5 | 3 | 53.9 | 3 |
| Gemini-2.5-Pro | 59.3 | 2 | 54.4 | 2 | 58.1 | 2 |
| GAR-8B | 62.2 | 1 | 57.4 | 1 | 61.8 | 1 |
| Model | GPT-4o | o3 | Gemini-2.5-Flash | Gemini-2.5-Pro | ||||
| Score | Rank | Score | Rank | Score | Rank | Score | Rank | |
| DAM-3B | 13.1 | 10 | 9.3 | 10 | 9.4 | 10 | 0.3 | 10 |
| PAM-3B | 21.1 | 9 | 21.5 | 9 | 25.3 | 9 | 31.3 | 9 |
| VP-SPHINX-13B | 32.3 | 8 | 27.5 | 8 | 30.4 | 8 | 32.7 | 8 |
| Qwen2.5-VL-32B | 36.8 | 7 | 28.1 | 7 | 37.4 | 7 | 41.8 | 7 |
| InternVL3-38B | 45.1 | 6 | 30.8 | 6 | 41.9 | 6 | 47.4 | 6 |
| GPT-4o | 51.5 | 5 | 31.8 | 5 | 46.3 | 5 | 50.6 | 5 |
| o3 | 56.9 | 4 | 37.8 | 3 | 52.8 | 4 | 60.7 | 3 |
| GAR-1B | 57.5 | 3 | 37.4 | 4 | 55.1 | 3 | 59.8 | 4 |
| Gemini-2.5-Pro | 59.3 | 2 | 44.5 | 2 | 57.4 | 2 | 61.3 | 2 |
| GAR-8B | 62.2 | 1 | 46.7 | 1 | 60.3 | 1 | 62.6 | 1 |
| Input Type | GPT-4o | o3 | Gemini-2.5-Pro | ||||||
| VQA | Cap | DLC-Bench | VQA | Cap | DLC-Bench | VQA | Cap | DLC-Bench | |
| Type 1 | 53.5 | 51.5 | 41.0 | 61.3 | 56.9 | 48.0 | 64.2 | 59.3 | 48.4 |
| Type 2 | 56.4 | 52.0 | 38.4 | 63.4 | 54.9 | 47.8 | 61.9 | 58.1 | 52.9 |
| Type 3 | 51.2 | 31.9 | 25.4 | 57.8 | 40.7 | 34.9 | 61.9 | 61.3 | 54.6 |
| Type 4 | 48.3 | 38.2 | 28.5 | 59.0 | 47.1 | 41.3 | 66.0 | 44.1 | 47.8 |
Robust Analysis of GAR-Bench. We conducted a subsampling stability analysis. We randomly subsampled GAR-Bench-VQA and GAR-Bench-Cap to 50% and 25% of their original sizes (for each subtask) and re-evaluated the full suite of models in Table 13 and Table 13, respectively. Our goal was to test whether the relative performance rankings remained consistent even with significantly fewer samples. The results, presented in the tables, demonstrate a high degree of ranking stability. As the results show, the relative ordering of models is remarkably stable. For example, in GAR-Bench-Cap, the top 3 models (GAR-8B, Gemini-2.5-Pro (DeepMind, 2025b), GAR-1B) and bottom 3 models (VP-SPHINX (Lin et al., 2025a), PAM (Lin et al., 2025b), DAM (Lian et al., 2025)) maintain their general ranking group even at a 1/4 sample size. While minor fluctuations exist among the top-tier models in the VQA 1/4 split (e.g., GAR-8B jumping from 3rd to 1st), the overall performance tiers are preserved.
Robust Analysis of LLM-Judges. To directly investigate the consistency and potential bias of the LLM judge, we conducted a new cross-judge validation experiment. We re-evaluated all models on GAR-Bench-Cap using four different powerful LLMs as judges: GPT-4o (OpenAI, 2024a) (our original judge), o3 (OpenAI, 2025), Gemini-2.5-Flash (DeepMind, 2025a), and Gemini-2.5-Pro (DeepMind, 2025b). The results, presented in Table 15, demonstrate a high degree of consistency in model rankings. While the absolute scores fluctuate across different judges, which reflects their inherent stylistic preferences or scoring strictness, the relative ranking of models is remarkably stable. Most importantly, our GAR-8B consistently ranks 1st across all four distinct judges, and our other GAR models also consistently place in the top tier. This experiment suggests that our primary claims about GAR’s superior performance are robust and not an artifact of a single judge’s bias.
Input Type Analysis for General Models. To investigate whether general models lack the ability to interpret masks or are genuinely deficient in understanding local details and relationships, we conduct an analysis using four different input types illustrated in Figure 7: (1) Type 1: Separated global image and local masks (our original setting). (2) Type 2: Drawing mask outlines directly onto the image. (3) Type 3: Using bounding boxes derived from masks. (4) Type 4: Using center-point crosses derived from masks. Empirical results are presented in Table 15, where powerful models consistently struggle across all four settings. This crucial finding reveals a fundamental deficiency in fine-grained perception and relational reasoning, regardless of how the regions of interest are indicated.
@@ D.1 Qualitative Results on GAR-Bench
“Relation” of GAR-Bench-VQA. In Figure 8, we provide qualitative comparisons on the “relation” protocol of our GAR-Bench-VQA, including two failure cases (the last row). As demonstrated in the figure, GAR-8B manages to not only effectively model relationships but also leverage crucial local details for choosing the best answer. For instance, in the right example of the middle row, the person (<Prompt0>) is actually not reading the book (<Prompt1>), since she is looking at the camera. Our GAR-8B manages to recognize such details and thus select “<Prompt0> is holding <Prompt1>” instead of “reading”, while both Gemini-2.5-Pro (DeepMind, 2025b) and o3 (OpenAI, 2025) fail.
However, as illustrated in the last two examples in Figure 8, current models still sometimes struggle to understand complex relationships with more than two objects. Constructing such complicated training data and keeping the correctness of relation annotations could be a potential solution.
“Non-Entity Recognition” of GAR-Bench-VQA. In Figure 9, we provide qualitative comparisons on the “non-entity recognition” protocol of our GAR-Bench-VQA, including two failure cases (the last row). As demonstrated in the figure, GAR-8B is able to correctly recognize objects in the mirror without any depth prior, thanks to its encoded global contexts.
However, as demonstrated in the right case in the last row, current models still struggle to distinguish whether the reflection actually comes from the mirror (<Prompt2>) or other reflective surfaces (<Prompt0> and <Prompt1>).